OpenStudy (anonymous):

Which equation represents a line perpendicular to the line whose equation is 2y − 5x = 10? A. 5y - 2x = 25 B. 5y + 2x = 10 C. 4y - 10x = 12 D. 2y + 10x = 8

OpenStudy (anonymous):

if it is perpendicular, then the slopes are opposite

OpenStudy (anonymous):

2y-5x=10 2y=5x+10 y=(5/2)x+5

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

you need to put all of your options into slope intercept form and whichever is (-2/5)x is it

OpenStudy (anonymous):

thx man